shift summary, day 22 June 99

Marj Corcoran (corcoran@physics.rice.edu)
Tue, 22 Jun 1999 16:07:42 -0500 (CDT)

Shift Summary for DAY shift, 22-June-99

On shift: A. Tripathi, M. Corcoran, C. Wolfe

Hours of beam delivered at all intensitites: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of E832 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time: Beam tuning, swap board in ktev2,
drt_fresh_start, TeV quench at end of shift

==================================================================

Runs ended this shift:

Run 13689: E832 data
Run 13690 E832 data
Run 13691 E832 data
Run 13692 E832 data
RUn 13693 E832 data

Tapes brought to FCC: None

==================================================================
Overview of shift: Mostly E832 data-taking. Some beam tuning, swap
of suspect board in ktev2

===================================================================

Description of shift:

Time Run Spill Comments
=======================================================================

7:00 Am 13689 Inherit run from owl. Intensity is
low, 2e12, problems with MI RF.

8:00 13689 Ed calls, offline analysis shows a
slight difference in intensity for the
two beams, Rick F may try to fix it.

9:30 13689 284 Rick F calls, he will be doing some
beam tuning. Stop run 13689.

10:25 13690 1 Rick F calls back, tuning done. Start
new run. Horizontal position on SEED2
is now +.2, vertical pos 6.2

Scaler window seems frozen, restart sdaq
That doesn't work, have to restart
ktscaler by hand.

10:55 13690 27 Accidently cycled power in laser crate
repeater, did kfrend and triggers
returned

13:30 13690 Ended run for swap of board on ktev2

At the same time we notice beam stop has
dropped in, MCR doesn't know why.

13:50 13691 1 KTeV2 back up, beam stop back out,
drt_fresh_start, new run.
Mouse on lower screen of atev2 is dead,
making it impossible to check things on
that sdaq screen

14:50 13691 36 Rick F calls to say beam is off
vertically due to NM2V not ramping.
NM2V was off for about 20 mins.
Beam was about 5 wires low.
Rick F has also tuned to
equalize intensity in E and W beams.
End run to do tuning.

14:50 13692 1 Restart data-taking.

15:40 13693 29 beam off, quench. Reboot atev2

===================================================================

Followup notes on previous shift:

===================================================================
Items to follow up on next shift:

Board in ktev2 was replaced this shift. Hopefully no more crashed of KTeV2.

Keep an eye on NM2V, it stopped ramping for no apparent reason.